# Theory of polarization-switchable electrical conductivity anisotropy in nonpolar semiconductors

### v1

@@ Abstract
Our analyses for the entire nonpolar crystallographic point groups are shown in Section B and Section C of our SM. This helps to identify a variety of point groups that enable the polarization-switchable ECA, as summarized in Table 2. The cases listed in Table 2 can be classified into Category A and Category B. Belonging to Category A are cases where for ECA between the $\bm{\chi}$ and $\bm{\chi}^{\prime}$ directions (driven by $P_{w}$ polarization) the point group symmetry protects the $\sigma_{\chi,\chi}(P_{w},\mu)-\sigma_{\chi^{\prime},\chi^{\prime}}(P_{w},\mu)=-\sigma_{\chi,\chi}(-P_{w},\mu)+\sigma_{\chi^{\prime},\chi^{\prime}}(-P_{w},\mu)$ relationship. In such sense, the polarization-switchable feature for the ECA is rigorous, being robust against higher-order corrections  [^5] By higher-order corrections, we mean corrections that go beyond effective mass approximation or involve higher-order contributions from electric polarization. Such higher-order corrections should be compatible with the corresponding point group symmetry.. For instance, the $\bar{4}3m$ point group contains two-fold rotation symmetry operations along the $\mathbf{x}$ and $\mathbf{y}$ directions. These symmetry operations protect the polarization-switchable feature for the $P_{z}$-induced ECA between the $\mathbf{x}+\mathbf{y}$ and $-\mathbf{x}+\mathbf{y}$ directions, because they provide a linkage between $\sigma_{xy,xy}(P_{z},\mu)-\sigma_{\bar{x}y,\bar{x}y}(P_{z},\mu)$ and $\sigma_{\bar{x}y,\bar{x}y}(-P_{z},\mu)-\sigma_{xy,xy}(-P_{z},\mu)$. The Category B contains cases for which the polarization-switchable ECA, derived within effective mass approximation (up to first-order in polarization), is not rigorously guaranteed by point group symmetry. As an example, the $\bar{6}2m$ point group hosts an ECA between the $\mathbf{x}$ and $\mathbf{y}$ directions driven by $P_{x}$ polarization, while this point group has no symmetry operations linking $P_{x}$ with $-P_{x}$ (see Section C of the SM). This means that the $\sigma_{x,x}(P_{x},\mu)-\sigma_{y,y}(P_{x},\mu)=\sigma_{y,y}(-P_{x},\mu)-\sigma_{x,x}(-P_{x},\mu)$ relationship associated with the $\bar{6}2m$ point group may be broken when involving higher-order corrections such as $\zeta_{xyz}^{lmn}P_{x}^{l}k_{x}^{m}k_{y}^{n}$ ($l,m,n$ being natural numbers with $l\geq 2$ or $m+n>2$).
@@
To end this section, we comment on the correlation between polarization-switchable ECA and polarization-reversal driven conductivity change. In Section A of the SM, we demonstrate that the $P_{w}$-switchable ECA between the $\bm{\chi}$ and $\bm{\chi^{\prime}}$ directions is rooted in $(\eta_{\chi\chi}+\kappa_{w\chi\chi}P_{w})k_{\chi}^{2}$ or $(\eta_{\chi\chi}-\kappa_{w\chi\chi}P_{w})k_{\chi^{\prime}}^{2}$. According to Eqs. (3)–(6), $(\eta_{\chi\chi}+\kappa_{w\chi\chi}P_{w})k_{\chi}^{2}$ and $(\eta_{\chi\chi}-\kappa_{w\chi\chi}P_{w})k_{\chi^{\prime}}^{2}$ terms imply the polarization-reversal driven conductivity changes along the $\bm{\chi}$ and $\bm{\chi^{\prime}}$ directions, respectively. As for Table 2, we conclude that every $(\chi,\chi^{\prime})$ entry below $P_{w}$ indicates the conductivity changes resulting from the reversal of $P_{w}$, that is, $\sigma_{\chi,\chi}(P_{w})-\sigma_{\chi,\chi}(-P_{w})\neq 0$ and $\sigma_{\chi^{\prime},\chi^{\prime}}(P_{w})-\sigma_{\chi^{\prime},\chi^{\prime}}(-P_{w})\neq 0$. The only exceptions are associated with $23$ and $\bar{4}3m$ point groups, where $\sigma_{w,w}(P_{w})-\sigma_{w,w}(-P_{w})=0$ holds for $w$ being $x$, $y$, or $z$.
@@
Figure 1: Schematizations of the crystal structures for $\bar{4}3m$ AlP (a) and $\bar{6}2m$ NaSrP (b). The $\mathbf{x}$, $\mathbf{y}$, and $\mathbf{z}$ unit vectors in both panels are perpendicular to each other. In panel (a), the $\mathbf{x^{\prime}}$ and $\mathbf{y^{\prime}}$ unit vectors are defined as $\mathbf{x^{\prime}}=(\mathbf{x}+\mathbf{y})/\sqrt{2}$ and $\mathbf{y^{\prime}}=(-\mathbf{x}+\mathbf{y})/\sqrt{2}$. The boxes enclosed by cyan solid lines (i.e., $xyz$ coordinate system) and black dashed lines (i.e., $x^{\prime}y^{\prime}z$ coordinate system) denote the unit cell of AlP and the super cell used in our simulations, respectively.
@@
Semiconductors with polarization-switchable ECA. We move on to verify our theory and identify semiconductors with polarization-switchable ECA, by electrical conductivity calculations based on first principles. As shown in Table 2, a variety of point groups host the ECA driven by polarization and rigorously switched by flipping the polarization. Of particular interest is the $\bar{4}3m$ point group, where a majority of technologically important semiconductors (i.e., zinc-blende type) belong to this point group Hinuma et al. (2014); Grüneis et al. (2014). In the absence of electric polarization, the electrical conductivities associated with the $\bar{4}3m$ point group is isotropic among the $\mathbf{x}$, $\mathbf{y}$, and $\mathbf{z}$ directions. When gaining a polarization along $\mathbf{z}$, the rigorously polarization-switchable ECA is enabled between the $\mathbf{x}+\mathbf{y}$ and $-\mathbf{x}+\mathbf{y}$ directions (see Table 2). To verify this, we take the zinc-blende type AlP semiconductor, with a band gap of $\sim$2.4 eV Richman (1968), as our platform. In Fig. 1(a), the box enclosed by cyan solid lines schematizes the unit cell of AlP (i.e., $xyz$ coordinate system), while the larger super cell (i.e., $x^{\prime}y^{\prime}z$ coordinate system) is the one employed in our numerical simulations. To polarize the AlP semiconductor, we apply external electric fields of ±$\pm 2$ and ±$\pm 4$ MV/cm along ±$\pm\mathbf{z}$ direction to such a material. This creates four different $P_{z}$ polarization states in AlP, and yields finite $(\sigma_{x^{\prime},x^{\prime}}-\sigma_{y^{\prime},y^{\prime}})/\tau$ conductivity differences being $P_{z}$ dependent [see Fig. 2(a)]. According to Fig. 1(a), such conductivity differences characterize the ECA between the $\mathbf{x}+\mathbf{y}$ and $-\mathbf{x}+\mathbf{y}$ directions. Remarkably, the conductivity differences associated with electric fields of $+4$ and $-4$ MV/cm ($+2$ and $-2$ MV/cm) are rather symmetric with respect to the horizontal axis. This confirms our theory that $P_{z}$ drives strictly polarization-switchable ECA in materials with $\bar{4}3m$ point group.
Table 2 also contains several cases associated with approximately polarization-switchable ECA. A representative case is the $\bar{6}2m$ point group which hosts ECA between the $\mathbf{x}$ and $\mathbf{y}$ directions — being driven by $P_{x}$ electric polarization but not rigorously switchable by reversing $P_{x}$. We confirm this by computing the electrical conductivities for the NaSrP semiconductor with $\bar{6}2m$ symmetry Dong and DiSalvo (2007). The crystal structure of NaSrP is schematized in Fig. 1(b). As shown in Fig. S2 of the SM, both the valence band maximum (VBM) and conduction band minimum (CBM) for NaSrP locate at the Brillouin Zone center. This implies that the $P_{x}$-dependent energy dispersion $\epsilon(\mathbf{k},P_{x})=\tau_{xx}(k_{x}^{2}+k_{y}^{2})+\tau_{zz}k_{z}^{2}+\lambda_{xxx}P_{x}(k_{x}^{2}-k_{y}^{2})$ [see Eq. S(49) of the SM] should be approximately valid for chemical potential $\mu$ being close to VBM or CBM. To create $P_{x}$ in NaSrP, we polarize it by applying external electric fields of ±$\pm 4$ MV/cm along ±$\pm\mathbf{x}$ direction. The resulting $(\sigma_{x,x}-\sigma_{y,y})/\tau$ conductivity differences, associated with electric fields of ±$\pm 4$ MV/cm, are basically symmetric with respect to the horizontal axis [see Fig. 2(b)]; The asymmetric feature becomes significant for $\mu$ being larger than 2.8 eV. This validates that the $\bar{6}2m$ point group enables approximately polarization-switchable ECA.
We further numerically check the polarization-reversal induced conductivity changes in AlP and NaSrP. As shown in Fig. S3 of the SM, the $+2$ MV/cm and $-2$ MV/cm electric fields yield diverse $\sigma_{x^{\prime},x^{\prime}}$ conductivities in AlP, which indicates conductivity changes induced by polarization reversal (associated with ±$\pm 2$ MV/cm electric fields). Consistent with Eq. (6), the conductivity changes associated with ±$\pm 4$ MV/cm electric fields are almost doubled compared with those associated with ±$\pm 2$ MV/cm electric fields. Similarly, the polarization-reversal induced conductivity changes are also predicted to occur in NaSrP (see Fig. S4 of the SM).
@@
Figure 2: The ECA in AlP and NaSrP as functions of chemical potential $\mu$, where $\mu=0$ is referred to as the corresponding valence band maximum. Panel (a): $(\sigma_{x^{\prime},x^{\prime}}-\sigma_{y^{\prime},y^{\prime}})/\tau$ in AlP driven by electric fields along ±$\pm\mathbf{z}$ direction. Panel (b): $(\sigma_{x,x}-\sigma_{y,y})/\tau$ in NaSrP driven by electric fields along ±$\pm\mathbf{x}$ direction. In panels (a) and (b), the unit for the vertical axes is $10^{19}/(\Omega\cdot m\cdot s)$.
Summary and outlook. Within effective mass approximation, we have developed a theory on polarization-switchable ECA in nonpolar semiconductors, where the ECA exhibits a first-order response to polarization and is reversible by flipping the polarization. We have shown that the polarization-switchable ECA is accompanied with conductivity changes driven by the reversal of the electric polarization. Hosted by a variety of point groups (summarized in Table 2), our discussed polarization-switchable ECA phenomena can be classified into two categories. Associated with the first category are cases for which the ECA is rigorously switchable by reversing the polarization. Such a polarization-switchable feature is protected by point group symmetry and is robust against higher-order corrections (e.g., beyond effective mass approximation). In contrast, the polarization-switchable feature for ECA belonging to the second category — valid in the regime of effective mass approximation and first-order response to polarization — may be broken when involving higher-order corrections. Our theory is further numerically validated by electrical conductivity calculations based on first principles. This yields the predictions of AlP and NaSrP as two representative semiconductors with polarization-switchable ECA and polarization-reversal driven conductivity change. We expect that our theory opens a route toward the design of next-generation electronic devices based on our aforementioned anisotropic charge transport and related phenomena.
@@
Acknowledgements. – We acknowledge the support from the National Natural Science Foundation of China (Grants Nos. 12274174, T2225013, 52288102, 52090024, and 12034009). L.B. thanks the Vannevar Bush Faculty Fellowship (VBFF) grant No. N00014-20-1-2834 from the Department of Defense and award No. DMR-1906383 from the National Science Foundation AMASE-i Program (MonArk NSF Quantum Foundry). H.J.Z. thanks the supports from “Xiaomi YoungScholar” Project, Arkansas High Performance Computing Center, and high-performance computing center of Jilin University.

### v2

@@ Abstract
Our analyses for the entire nonpolar crystallographic point groups are shown in Section B and Section C of our SM. This helps to identify a variety of point groups that enable the polarization-switchable ECA, as summarized in Table 2. These point groups are associated with nonpolar piezoelectric materials. The cases listed in Table 2 can be classified into Category A and Category B. Belonging to Category A are cases where for ECA between the $\bm{\chi}$ and $\bm{\chi}^{\prime}$ directions (driven by $P_{w}$ polarization) the point group symmetry protects the $\sigma_{\chi,\chi}(P_{w},\mu)-\sigma_{\chi^{\prime},\chi^{\prime}}(P_{w},\mu)=-\sigma_{\chi,\chi}(-P_{w},\mu)+\sigma_{\chi^{\prime},\chi^{\prime}}(-P_{w},\mu)$ relationship. In such sense, the polarization-switchable feature for the ECA is rigorous, being robust against higher-order corrections  [^6] By higher-order corrections, we mean corrections that go beyond effective mass approximation or involve higher-order contributions from electric polarization. Such higher-order corrections should be compatible with the corresponding point group symmetry.. For instance, the $\bar{4}3m$ point group contains two-fold rotation symmetry operations along the $\mathbf{x}$ and $\mathbf{y}$ directions. These symmetry operations protect the polarization-switchable feature for the $P_{z}$-induced ECA between the $\mathbf{x}+\mathbf{y}$ and $\mathbf{x}+\mathbf{y}$ directions, because they provide a linkage between $\sigma_{xy,xy}(P_{z},\mu)-\sigma_{\bar{x}y,\bar{x}y}(P_{z},\mu)$ and $\sigma_{\bar{x}y,\bar{x}y}(-P_{z},\mu)-\sigma_{xy,xy}(-P_{z},\mu)$. The Category B contains cases for which the polarization-switchable ECA, derived within effective mass approximation (up to first-order in polarization), is not rigorously guaranteed by point group symmetry. As an example, the $\bar{6}2m$ point group hosts an ECA between the $\mathbf{x}$ and $\mathbf{y}$ directions driven by $P_{x}$ polarization, while this point group has no symmetry operations linking $P_{x}$ with $P_{x}$ (see Section C of the SM). This means that the $\sigma_{x,x}(P_{x},\mu)-\sigma_{y,y}(P_{x},\mu)=\sigma_{y,y}(-P_{x},\mu)-\sigma_{x,x}(-P_{x},\mu)$ relationship associated with the $\bar{6}2m$ point group may be broken when involving higher-order corrections such as $\zeta_{xyz}^{lmn}P_{x}^{l}k_{x}^{m}k_{y}^{n}$ ($l,m,n$ being natural numbers with $l\geq 2$ or $m+n>2$).
Our aforementioned polarization-switchable ECAs in nonpolar piezoelectrics are volatile, that is, an electric field is required to maintain the ECAs. A ferroelectric has a spontaneous electric polarization, and the reversal of the polarization involves an intermediate phase. If the point group of this intermediate phase belongs to the cases listed in Table 2, the ferroelectric may host ferroelectrically switchable ECA (i.e., nonvolatile). For instance, a nonpolar phase with $\bar{4}m2$ point group showcases $P_{z}$-switchable ECA between the $\mathbf{x}$ and $\mathbf{y}$ directions (see Table 2). By gaining a spontaneous $P_{z}$ polarization, this phase becomes ferroelectric with the $mm2$ point group, and such a ferroelectric phase enables ferroelectrically switchable $\sigma_{x,x}(P_{z},\mu)-\sigma_{y,y}(P_{z},\mu)$ anisotropic conductivity. Following this logic, we derive guidelines for screening ferroelectrics with ferroelectrically switchable ECA (summarized in Table 3).
Table 3: Ferroelectrically switchable ECA. The first column lists the point groups for intermediate phases. Gaining spontaneous polarizations drive these phases ferroelectric, and the resulting polar point groups are listed in the second column; Each spontaneous polarization is shown as $P_{x}$, $P_{y}$, or $P_{z}$ in the parenthesis. The $(\chi,\chi^{\prime})$ in the third column indicates the ferroelectrically switchable ECA between $\bm{\chi}$ and $\bm{\chi^{\prime}}$. The subscripts (if existing) in the polar point group symbols indicate the directions of the associated symmetry operations. To be specific, $\alpha$ in $2_{\alpha}$ labels the direction for the twofold rotation axis, and $\beta$ in $m_{\beta}$ marks the normal direction for the mirror plane ($\alpha$ and $\beta$ being $x$, $y$, $z$, $\cdots$).
Category A: rigorous cases
Intermediate Ferroelectric ECA
$\bar{4}$ $2_{z}$ ($P_{z}$) $(x,y)$; $(xy,\bar{x}y)$
$\bar{4}2m$ $m_{xy}m_{x\bar{y}}2_{z}$ ($P_{z}$) $(xy,\bar{x}y)$
$\bar{4}m2$ $m_{x}m_{y}2_{z}$ ($P_{z}$) $(x,y)$
$312$ 11 ($P_{x}$) $(xy,\bar{x}y)$
$321$ 11 ($P_{y}$) $(xy,\bar{x}y)$
$\bar{6}m2$ $m_{z}$ ($P_{x}$) $(xy,\bar{x}y)$
$\bar{6}2m$ $m_{z}$ ($P_{y}$) $(xy,\bar{x}y)$
$23$ $2_{x}$ ($P_{x}$) $(yz,\bar{y}z)$
$23$ $2_{y}$ ($P_{y}$) $(zx,\bar{z}x)$
$23$ $2_{z}$ ($P_{z}$) $(xy,\bar{x}y)$
$\bar{4}3m$ $m_{yz}m_{y\bar{z}}2_{x}$ ($P_{x}$) $(yz,\bar{y}z)$
$\bar{4}3m$ $m_{zx}m_{z\bar{x}}2_{y}$ ($P_{y}$) $(zx,\bar{z}x)$
$\bar{4}3m$ $m_{xy}m_{x\bar{y}}2_{z}$ ($P_{z}$) $(xy,\bar{x}y)$
Category B: approximate cases
Intermediate Ferroelectric ECA
$\bar{6}$ $m_{z}$ ($P_{x}$) $(x,y)$; $(xy,\bar{x}y)$
$\bar{6}$ $m_{z}$ ($P_{y}$) $(x,y)$; $(xy,\bar{x}y)$
$\bar{6}m2$ $m_{z}m_{x}2_{y}$ ($P_{y}$) $(x,y)$
$\bar{6}2m$ $m_{y}m_{z}2_{x}$ ($P_{x}$) $(x,y)$
$23$ $2_{x}$ ($P_{x}$) $(\bar{y}z,x)$; $(x,yz)$
$23$ $2_{y}$ ($P_{y}$) $(\bar{z}x,y)$; $(y,zx)$
$23$ $2_{z}$ ($P_{z}$) $(\bar{x}y,z)$; $(z,xy)$
$\bar{4}3m$ $m_{yz}m_{y\bar{z}}2_{x}$ ($P_{x}$) $(\bar{y}z,x)$; $(x,yz)$
$\bar{4}3m$ $m_{zx}m_{z\bar{x}}2_{y}$ ($P_{y}$) $(\bar{z}x,y)$; $(y,zx)$
$\bar{4}3m$ $m_{xy}m_{x\bar{y}}2_{z}$ ($P_{z}$) $(\bar{x}y,z)$; $(z,xy)$
@@
To end this section, we comment on the correlation between polarization-switchable ECA and polarization-reversal driven conductivity change. In Section A of the SM, we show that the $P_{w}$-switchable ECA between the $\bm{\chi}$ and $\bm{\chi^{\prime}}$ directions is rooted in $(\eta_{\chi\chi}+\kappa_{w\chi\chi}P_{w})k_{\chi}^{2}$ or $(\eta_{\chi\chi}-\kappa_{w\chi\chi}P_{w})k_{\chi^{\prime}}^{2}$ term. According to Eqs. (3)–(6), $(\eta_{\chi\chi}+\kappa_{w\chi\chi}P_{w})k_{\chi}^{2}$ and $(\eta_{\chi\chi}-\kappa_{w\chi\chi}P_{w})k_{\chi^{\prime}}^{2}$ terms imply the polarization-reversal driven conductivity changes along the $\bm{\chi}$ and $\bm{\chi^{\prime}}$ directions, respectively. As for Table 2, we conclude that every $(\chi,\chi^{\prime})$ entry below $P_{w}$ indicates the conductivity changes resulting from the reversal of $P_{w}$, that is, $\sigma_{\chi,\chi}(P_{w},\mu)-\sigma_{\chi,\chi}(-P_{w},\mu)\neq 0$ and $\sigma_{\chi^{\prime},\chi^{\prime}}(P_{w},\mu)-\sigma_{\chi^{\prime},\chi^{\prime}}(-P_{w},\mu)\neq 0$. The only exceptions are associated with $23$ and $\bar{4}3m$ point groups, where $\sigma_{w,w}(P_{w},\mu)-\sigma_{w,w}(-P_{w},\mu)=0$ holds for $w$ being $x$, $y$, or $z$. These arguments are readily extended to the cases shown in Table 3.
@@
Figure 1: Schematizations of the crystal structures for $\bar{4}3m$ AlP (a) and $mm2$ KH2PO4 (b). The $\mathbf{x}$, $\mathbf{y}$, and $\mathbf{z}$ unit vectors in both panels are perpendicular to each other $\mathbf{z}$ in panel (b) being omitted]. In panel (a), $\mathbf{x^{\prime}}$ and $\mathbf{y^{\prime}}$ are defined as $\mathbf{x^{\prime}}=(\mathbf{x}+\mathbf{y})/\sqrt{2}$ and $\mathbf{y^{\prime}}=(-\mathbf{x}+\mathbf{y})/\sqrt{2}$. The boxes enclosed by cyan solid lines (i.e., xyzxyz coordinate system) and black dashed lines (i.e., $x^{\prime}y^{\prime}z$ coordinate system) denote the unit cell of AlP and the super cell used in our simulations, respectively. In panel (b), the K ions are not shown.
III.3 C. Semiconductors with polarization-switchable ECA
We move on to identify semiconductors with polarization-switchable ECA, by electrical conductivity calculations based on first principles. As shown in Tables 2 and 3, a variety of point groups host polarization-dependent ECAs that are rigorously switched by flipping the polarization  [^7] Tables 2 and 3 also contain cases associated with approximately polarization-switchable ECA. We demonstrate and verify this by computing the electrical conductivities for the NaSrP semiconductor, which is shown in Section D of our SM.. We select $\bar{4}3m$ in Table 2 and $\bar{4}m2$ in Table 3 as our platforms, where the former is the point group of various zinc-blende type semiconductors (e.g., AlP [64]) being technologically important [65, 66] and the latter is the point group of the intermediate paraelectric phase for ferroelectric KH2PO4 [67, 68, 69].
@@
We use AlP and KH2PO4 to validate our theory. In Fig. 1(a), the box enclosed by cyan solid lines schematizes the unit cell of AlP (i.e., xyzxyz coordinate system), while the larger super cell (i.e., $x^{\prime}y^{\prime}z$ coordinate system) is the one employed in our numerical simulations. To polarize the AlP semiconductor, we apply external electric fields of $\pm 2$ and $\pm 4$ MV/cm along $\mathbf{z}$ direction to such a material. This creates four different $P_{z}$ polarization states in AlP, and yields finite $(\sigma_{x^{\prime},x^{\prime}}-\sigma_{y^{\prime},y^{\prime}})$ conductivity differences being $P_{z}$ dependent [see Fig. 2(a)]. According to Fig. 1(a), such conductivity differences characterize the ECA between the $\mathbf{x}+\mathbf{y}$ and $\mathbf{x}+\mathbf{y}$ directions. The conductivity differences associated with electric fields of +4+4 and 4-4 MV/cm (+2+2 and 2-2 MV/cm) are rather symmetric with respect to the horizontal axis at $(\sigma_{x^{\prime},x^{\prime}}-\sigma_{y^{\prime},y^{\prime}})=0$. This confirms our theory that $P_{z}$ drives strictly polarization-switchable ECA in materials with $\bar{4}3m$ point group. Fig. 2(b) shows $\Delta\sigma=|\sigma_{x^{\prime},x^{\prime}}-\sigma_{y^{\prime},y^{\prime}}|$ and $\sigma_{0}=(\sigma_{x^{\prime},x^{\prime}}+\sigma_{y^{\prime},y^{\prime}})/2$ for AlP, associated with an electric field of +4+4 MV/cm along $\mathbf{z}$. Strikingly, $\Delta\sigma$ is sizable as compared to $\sigma_{0}$, in the vicinity of valence band maximum. We use $\Delta\sigma/\sigma_{0}$ to describe the relative anisotropic conductivity, and this resembles the approach for quantifying the relative anisotropic magnetoresistance (see e.g., Ref. [16]). The $\Delta\sigma/\sigma_{0}$ ranges from $\sim 125\%$ to $\sim 75\%$ when decreasing the chemical potential $\mu$ from $0.0$ to 0.1-0.1 eV. Experimentally, such a variation of $\mu$ is achievable by doping AlP with Be which creates $p$-type hole carriers [70].
Ferroelectric KH2PO4 has a point group of $mm2$ [67, 68, 69] and a spontaneous polarization along $\mathbf{z}$ [see Fig. 1(b)]. The point group of the intermediate paraelectric phase, for polarization reversal, is $\bar{4}m2$ [67, 68, 69]. According to Table 3, KH2PO4 exhibits ferroelectrically switchable anisotropic conductivity between $\mathbf{x}$ and $\mathbf{y}$, which stems from spontaneous polarization and hence are nonvolatile. Our assertions are validated by Fig. 3(a). As shown in Fig. 3(b), $\Delta\sigma=|\sigma_{x,x}-\sigma_{y,y}|$ is sizable compared with $\sigma_{0}=(\sigma_{x,x}+\sigma_{y,y})/2$. In the range of $0.1~\mathrm{eV}\leq\mu\leq 0~\mathrm{eV}$, the relative anisotropic conductivity $\Delta\sigma/\sigma_{0}$ varies between $50\%$ and $76\%$.
@@
Figure 2: The ECA in AlP as functions of chemical potential $\mu$, where $\mu=0$ is referred to as the corresponding valence band maximum. Panel (a): $(\sigma_{x^{\prime},x^{\prime}}-\sigma_{y^{\prime},y^{\prime}})$ driven by electric fields along $\pm\mathbf{z}$ direction. Panel (b): $(\sigma_{x^{\prime},x^{\prime}}+\sigma_{y^{\prime},y^{\prime}})/2$ and $|\sigma_{x^{\prime},x^{\prime}}-\sigma_{y^{\prime},y^{\prime}}|$ associated with an electric field of +4+4 MV/cm along the $\mathbf{z}$ direction. The unit for the vertical axes is $10^{19}\tau/(\Omega\cdot m\cdot s)$.
We now check the conductivity changes in AlP and KH2PO4 that are induced by the reversal of polarization. We polarize AlP by applying $\pm 4$ MV/cm electric fields along the $\mathbf{z}$ direction. As shown in Fig. S5 of the SM, this yields $\pm P_{z}$ polarization states with unequal $\sigma_{1}=\sigma_{x^{\prime},x^{\prime}}(P_{z},\mu)$ and $\sigma_{2}=\sigma_{x^{\prime},x^{\prime}}(-P_{z},\mu)$ conductivities. Following the definition of tunneling magnetoresistance ratio [71], we define our conductivity change ratio as $|\sigma_{1}-\sigma_{2}|/\sigma_{\mathrm{min}}$, where $\sigma_{\mathrm{min}}$ is the minimum between $\sigma_{1}$ and $\sigma_{2}$. At $\mu=0~\mathrm{eV}$, the conductivity change ratio in AlP reaches $\sim 333\%$. As for KH2PO4, switching the direction of its polarization between $\mathbf{z}$ and $\mathbf{z}$ yields diverse $\sigma_{1}=\sigma_{x,x}(P_{z},\mu)$ and $\sigma_{2}=\sigma_{x,x}(-P_{z},\mu)$ conductivities (see Fig. S6 of the SM), with a conductivity change ratio of $\sim 104\%$ at $\mu=0~\mathrm{eV}$.
Figure 3: The ECA in ferroelectric KH2PO4 as functions of chemical potential $\mu$, where $\mu=0$ is referred to as the corresponding valence band maximum. Panel (a): $(\sigma_{x,x}-\sigma_{y,y})$ associated with electric polarizations along $\pm\mathbf{z}$ direction. Panel (b): $(\sigma_{x,x}+\sigma_{y,y})/2$ and $|\sigma_{x,x}-\sigma_{y,y}|$ associated with an electric polarization along the $\mathbf{z}$ direction. The unit for the vertical axes is $10^{19}\tau/(\Omega\cdot m\cdot s)$.
IV IV. Summary and outlook
Within effective mass approximation, we have developed a theory on polarization-switchable ECA in semiconductors, where the ECA exhibits a first-order response to electric polarization and is reversible by flipping the polarization. We have shown that the polarization-switchable ECA is accompanied with polarization-reversal induced conductivity change. This may occur in piezoelectrics or ferroelectrics with their possible crystallographic point groups summarized in Tables 2 and 3. Our theory is validated by electrical conductivity calculations based on first principles. This yields the predictions of piezoelectric AlP and ferroelectric KH2PO4 as two representative semiconductors with the aforementioned transport phenomena.
In the vicinity of the valence band maximum, our estimated relative anisotropic conductivities in AlP (driven by $\pm 4$ MV/cm electric fields) and KH2PO4 (occurred spontaneously) exceed $50\%$. Such values are not less than the relative anisotropic magnetoresistance in permalloy (e.g., between $16\%$ and $25\%$ for Ni80Fe20 [16]) that is used in industry. At $\mu=0$ eV, AlP and KH2PO4 exhibit polarization-reversal induced conductivity changes with change ratios of $\sim 333\%$ and $\sim 104\%$, respectively. The ratio in KH2PO4 is lower than the magnetoresistance ratio in commercial magnetic tunnel junctions (usually, larger than $250\%$[72], while the ratio in AlP exceeds $250\%$. In view of this, polarization-switchable electrical conductivity anisotropy and polarization-reversal induced conductivity change in AlP should be promising for device design. As an outlook, we expect that our work opens a route toward next-generation electronic devices based on anisotropic charge transport and related phenomena. For sake of nonvolatile devices, the discovery of advanced ferroelectrics with ferroelectrically switchable anisotropic conductivity and ferroelectric-reversal induced conductivity change (being sizable) is of particular interest.
@@
Acknowledgements. We acknowledge the support from the National Natural Science Foundation of China (Grants Nos. 12274174, T2225013, 52288102, 52090024, and 12034009). L.B. thanks the Vannevar Bush Faculty Fellowship (VBFF) grant No. N00014-20-1-2834 from the Department of Defense and award No. DMR-1906383 from the National Science Foundation AMASE-i Program (MonArk NSF Quantum Foundry). H.J.Z. thanks the supports from “Xiaomi YoungScholar” Project, Arkansas High Performance Computing Center, and high-performance computing center of Jilin University. The authors are grateful to Prof. Zheng Wen at Qingdao University for the valuable discussion.